Coronado Industries had January 1 inventory of $300000 when it adopted dollar-value LIFO. During the year, purchases were $1830000 and sales were $3070000. December 31 inventory at year-end prices was $388640, and the price index was 112. What is Coronado Industries’s gross profit? please show work

a. $1292640.

b. $1328640.

c. $1275000.

d. $2717360.
Answer #1

Calculation of COGS

$388,640 ÷ 1.12 = $347,000 – $300,000 = $47,000.

300,000 + (47,000 x 1.12) = 352,640

COGS = 300,000 + 1,830,000 - 352,640

=1,777,360

Gross Profit = 3,070,000 - 1,777,360

=1,292,640
Answer #2

To calculate the gross profit using dollar-value LIFO, we need to follow these steps:

Step 1: Calculate the ending inventory at base-year prices. Step 2: Calculate the cost of goods sold (COGS) at base-year prices. Step 3: Calculate the gross profit.

Given data: January 1 inventory (base-year prices) = $300,000 Purchases (base-year prices) = $1,830,000 Sales (base-year prices) = $3,070,000 December 31 inventory (year-end prices) = $388,640 Price index = 112

Step 1: Calculate the ending inventory at base-year prices. Ending inventory at year-end prices = $388,640 Ending inventory at base-year prices = Ending inventory at year-end prices / Price index Ending inventory at base-year prices = $388,640 / 1.12 Ending inventory at base-year prices = $347,000

Step 2: Calculate the cost of goods sold (COGS) at base-year prices. COGS at base-year prices = Beginning inventory at base-year prices + Purchases at base-year prices - Ending inventory at base-year prices COGS at base-year prices = $300,000 + $1,830,000 - $347,000 COGS at base-year prices = $1,783,000

Step 3: Calculate the gross profit. Gross profit = Sales (base-year prices) - COGS (base-year prices) Gross profit = $3,070,000 - $1,783,000 Gross profit = $1,287,000

So, the correct answer is: a. $1,287,000.

answered by: Hydra Master